OpenStudy (anonymous):

Which equation has no solution? A)8 – (5v + 3) = 5v – 5 B)3m – 6 = 5m + 7 – m C)3w + 4 – w = 5w – 2(w – 2) D)7y + 9 = 7y – 6

OpenStudy (anonymous):

If there is only one, then it is D) 7y + 9 = 7Y -6

OpenStudy (anonymous):

7y + 9 = 7y – 6 subtract 7y from both sides 7y - 7y + 9 = 7y - 7y - 6 9 = -6 this is absurd as 9 can never be equal to -6 so this equation has no solution